# Collective optical properties of moiré excitons

### v1

@@ Abstract
We propose that excitons in moiré transition metal dichalcogenide bilayers offer a promising platform for investigating collective radiative properties. While some of these optical properties resemble those of cold atom arrays, moiré excitons extend to the deep subwavelength limit, beyond the reach of current optical lattice experiments. Remarkably, we show that the collective optical properties can be exploited to probe certain correlated electron states. Specifically, we illustrate that the Wigner crystal states of electrons doped into these bilayers act as an emergent periodic potential for excitons. Moreover, the collective dissipative excitonic bands and their associated Berry curvature can reveal various charge orders that emerge at the corresponding electronic doping. Our study provides a promising pathway for future research on the interplay between collective effects and strong correlations involving moiré excitons.
@@
On the other hand, excitons (electron-hole bound states) in transition metal dichalcogenide (TMD) bilayers [6, 7, 8, 9, 10, 11] can form in a moiré lattice, providing an interesting platform to study collective optical properties. Moreover, the bilayers can host Wigner crystal (WC) states of doped electrons  [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], which can modify the excitonic properties.
In this work, we investigate the collective optical properties of moiré excitons in TMDs at incompressible states of doped electrons for various fractional fillings $\nu_{e}$. In these cases, the localized charges provide an additional superpotential, which effectively modifies the lattice for excitons (see Fig. 1) and their collective optical properties. Consequently, distinct spectral and topological properties of the excitons emerge at different WCs, offering optical signatures of these charge orders (see Table 1). Notably, such filling fractions can be easily accessed in gate-tunable experiments. Such tunability, combined with adjustable twist angles, further allows for the exploration of subwavelength emitter-array physics within different lattice configurations.
Figure 1: (a) Illustration of tunneling driven by dipole-dipole interaction (green arrow) of an exciton, as indicated by a pair of red and light blue dots, in a moiré potential (yellow) with period $a_{M}$ in the absence of doped electrons ($\nu_{e}=0$). (b) At certain finite dopings (here $\nu_{e}=\frac{1}{3}$), doped electrons (dark blue dots) form a Wigner crystal state. Due to exciton-electron repulsion, the exciton cannot populate the sites occupied by electrons (indicated by the crossed-out path). The remaining available sites form the emergent lattice.
@@
Two-band model. — We consider the lowest conduction (CB) and the highest valence (VB) bands of two TMD monolayers. Each band is labeled by the valley pseudospin, which is locked to the real spin degrees of freedom for energy scales lower than the corresponding spin-orbit splittings [26]. Stacking these layers with a tunable twist angle or (for heterobilayers) lattice mismatch generally leads to interlayer coupling with an enlarged spatial periodicity $a_{M}$. We specifically consider heterobilayers [27], as the band offsets therein allow us to capture the effect of interlayer tunneling by emergent superlattice potentials, which split the original electronic dispersion into moiré bands  [^2] Such band misalignment also suppresses the topological effects favored by layer-degenerate states [67], which is not the main focus of our work.. In addition, we consider doping electrons into the first conduction moiré band and focus on the regime with dilute electron-hole pairs [6, 29], generated via weak optical excitation. Note that within such a dilute regime, the emission properties are captured by the single excitation subspace, which can be described by a non-Hermitian Hamiltonian  [^3] In this case, it is equivalent to the full Lindbladian dynamics. While our formalism can be straightforwardly generalized, for simplicity, all (doped and optically excited) electrons are assumed to be in the same layer.
@@
Zero doping. — At $\nu_{e}=0$, optical excitations from Eq. (1) are given by excitons. To capture the lowest composite particle, we employ the tight-binding approximation to the two-particle Hamiltonian operator from $\hat{\mathcal{H}}$, whose eigenfunctions in the tight-binding limit can be approximated by moiré-Wannier orbitals [31]. The lowest energy orbital at each supersite $\bm{R}$, denoted as $w_{\bm{R}}(\bm{s}_{c},\bm{s}_{v})$ with $\bm{s}_{c}$ and $\bm{s}_{v}$ being the coordinates of the electron and hole, respectively, defines the corresponding exciton creation operator:
@@
Projecting $\hat{\mathcal{H}}$ onto these basis states, the superlattice Hamiltonian becomes  [^8] Here, we drop the dipole-dipole interaction from the small out-of-plane transition dipole (not the permanent dipole) of interlayer excitons [43], which can be further suppressed via hybridization with intralayer excitons [69].:
@@
Finite doping with charge order. — We can generalize Eq. (4) to finite $\nu_{e}$, where doped electrons form WCs. These charge orders could emerge from a generalized Hubbard model [44, 45], which naturally appears by projecting Eq. (1) onto the first moiré band of doped charges [31]. Assuming that the electronic states being stable against the dynamics of dilute optical excitations, we can treat the WC as a spatially periodic detuning with strength characterized by $\omega_{t}-\omega_{\mathrm{ex}}$, where $\omega_{t}$ is the energy of the lowest three-body state (one doped electron together with the excited electron-hole pair) within a supercell. In particular, in the regime $|\omega_{t}-\omega_{\mathrm{ex}}|\gg|t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}|$, the low-energy optical excitations are either within the $\sim\omega_{\mathrm{ex}}$ or $\sim\omega_{t}$ manifold, depending on the sign of $\omega_{t}-\omega_{\mathrm{ex}}$. We focus specifically on the case $\omega_{\mathrm{ex}}\ll\omega_{t}$, which is realizable by interlayer excitons [31], to study the collective behavior of these emitters. In this subspace, these excitons are still described by the Hamiltonian Eq. (4101010Note that the energy integrals are different for the excitations in the $\sim\omega_{\mathrm{ex}}$ and $\sim\omega_{t}$ manifolds, due to their distinction in wavefunctions, except that now $\{\bm{R}\}$ lies in the emergent lattices, which are complementary lattices of the WCs. Accordingly, different emergent lattices can be realized simply by accessing distinct WCs in moiré TMD bilayers, which can be tuned by the electron filling fraction $\nu_{e}$ (doped electrons per supercell) [44, 14, 25]. We list a few possible emergent lattices observed in WSe2/WS2 in Table 1.
@@
Figure 2: Collective excitonic lineshifts $\Delta_{\bm{Q}}$ (solid lines) and linewidths $\Gamma_{\bm{Q}}$ (dashed) emerging from charge-ordered zero-twist WSe2/WS2 with electron fillings $\nu_{e}=0$ (a) and $\nu_{e}=\frac{1}{3}$ (b). The vertical axes are displayed in units of $\gamma N_{\nu_{e}}$, with $N_{\nu_{e}}$ defined in Eq. (6). The horizontal axes show Bloch momenta $\bm{Q}$ at high symmetry points, which follow a piecewise-linear path in the Brillouin zone, as depicted by the dashed hexagon in the inset of (a). Momenta within the light cone are indicated by the green shaded area (size enlarged for clarity). Different colors label distinct single-particle exciton bands. The parameters used are: $a_{M}=8.25$nm, $\omega_{\mathrm{ex}}=1.55$eV, and $a_{W}=2$nm [31, 18].
@@
Collective bands. — We proceed to study the Liouvillian spectrum incorporating tunneling driven by coherent and incoherent dipole-dipole interactions. To capture the relevant physics, we set $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}=0$ in Eq. (5) throughout this work  [^11] Neglecting $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}$ is a valid approximation at large $a_{M}$ because it decays exponentially with $|\bm{R}-\bm{R}|/a_{M}$, whereas the dipole-dipole interaction only scales as a power law.. Diagonalization of the Hamiltonian Eq. (4) yields the emitter spectrum $\omega_{\mathrm{ex}}+\Delta_{\bm{Q}}-\frac{i}{2}\Gamma_{\bm{Q}}$ characterized by the center-of-mass Bloch momentum $\bm{Q}$. Note that, if the excitonic states at all $\bm{R}$ constructively interfere, $\Gamma_{\bm{Q}}$ should scale with $\gamma N_{\nu_{e}}(\theta)$ [1], where:
$N_{\nu_{e}}(\theta)=(1+\theta^{2}/\delta^{2})N_{\nu_{e}},\quad N_{\nu_{e}}=\lambda_{\mathrm{ex}}^{2}/\mathcal{A}_{\nu_{e}},$ (6)
@@
with $\mathcal{A}_{\nu_{e}}$ denoting the emergent unit cell area at electron filling $\nu_{e}$ and zero twist, $\delta$ defining the lattice mismatch between the two monolayers [18]. As such, we present the collective spectrum in units of $\gamma N_{\nu_{e}}(\theta)$ hereafter to indicate the extent of constructive interference of the exciton eigenstates.
@@
Aside from these common properties, collective bands exhibit several distinct features at different $\nu_{e}$. We specifically illustrate how $\Delta_{0}$ and $\Gamma_{0}$ vary with $\nu_{e}$ in Fig. 3(a), as these quantities are directly accessible via reflection measurements at normal incidence [1]. First, different emergent lattices provide a distinct number of levels  [^12] Although the number of collective bands should equal two (coupled valleys) times the number of sublattices per unit emergent cell, we note that some states are degenerate and non-resolvable at high symmetry momenta.. Although this serves as the most direct signature to distinguish emergent lattices at each $\nu_{e}$, we note that states other than the lowest one possess a suppressed decay rate (except at $\nu_{e}=\frac{1}{4}$ and $\frac{1}{2}$), such that we expect they barely contribute to optical measurements. In contrast, the lowest state at each $\nu_{e}$, which is typically the most radiative one, could respond nontrivially to optical probe light; their collective properties — the overall increasing $\Delta_{0}$ and decreasing $\Gamma_{0}$ with $\nu_{e}$ — could serve as characteristics of the corresponding emergent lattices.
@@
Finally, these spectral properties can also be tuned via twisting angle $\theta$. Fig. 3(b) illustrates the dependence of $\Delta_{0}$ and $\Gamma_{0}$ on $N_{\nu_{e}}(\theta)$. Crucially, for the same type of lattice geometry, the collective linewidths scale inversely with the emergent unit cell area, which can be confirmed by the analytic expression of the Green’s tensor at $\bm{Q}=0$. These scalings provide an indirect probe for the emergent lattice geometry: for instance, all triangular arrays (e.g., $\nu_{e}=0,\frac{2}{3}$) belong to the same line in Fig. 3(b).

### v2

@@ Abstract
We propose that excitons in moiré transition metal dichalcogenide bilayers offer a promising platform for investigating collective radiative properties. While some of these optical properties resemble those of cold atom arrays, moiré excitons extend to the deep subwavelength limit, beyond the reach of current optical lattice experiments. Remarkably, we show that the collective optical properties can be exploited to probe certain correlated electron states without requiring subwavelength spatial resolution. Specifically, we illustrate that the Wigner crystal states of electrons doped into these bilayers act as an emergent periodic potential for excitons. Moreover, the collective dissipative excitonic bands and their associated Berry curvature can reveal various charge orders that emerge at the corresponding electronic doping. Our study provides a promising pathway for future research on the interplay between collective effects and strong correlations involving moiré excitons.
@@
On the other hand, excitons (electron-hole bound states) in transition metal dichalcogenide (TMD) bilayers [6, 7, 8, 9, 10, 11] can form in a moiré lattice, providing an interesting platform to study collective optical properties. Moreover, the bilayers can host Wigner crystal (WC) states of doped electrons  [12, 13, 14, 15, 16, 17, 18, 19, 20], which can modify the excitonic properties.
In this work, we investigate the collective optical properties of moiré excitons in TMDs at incompressible states of doped electrons for various fractional fillings $\nu_{e}$  [^1] For clarity, the term “collective” in this work refers to the interference effects of states of an optical emitter from different sites [22, 23] rather than many-body physics of multiple emitters [24, 25, 26]. We focus on a weak excitation regime that yields dilute excitons, allowing us to neglect their electrostatic repulsion that manifests in high-power experiments [26, 20, 16, 18]. These excitations tunnel between empty moiré sites through dipole-dipole interaction, but cannot sit on top of a doped electron due to exciton-electron repulsion, as illustrated in Fig. 1(a). Therefore, WCs emerging at finite $\nu_{e}$ act as additional superpotentials, modifying the lattice potential experienced by a single exciton. Moreover, different $\nu_{e}$’s give rise to distinct WCs [see Fig. 1(b)], which induce distinctive excitonic dynamics with quantitatively different collective (spectral and topological) excitonic properties [27]. As such, those optical signatures serve as fingerprints of the corresponding WCs (see Table 1). Notably, the filling fractions $\nu_{e}$ can be easily accessed in gate-tunable experiments. Such tunability, combined with adjustable twist angles, further allows for the exploration of subwavelength emitter-array physics within different lattice configurations.
Figure 1: (a) Illustration of tunneling mediated by dipole-dipole interactions (green arrow) of an exciton (indicated by a pair of red and light blue dots) in a moiré potential (yellow) with period $a_{M}$. The potential can host a Wigner crystal state (WC) of doped electrons (dark blue dots), here a filling of $\nu_{e}=\frac{1}{3}$ is illustrated. Due to exciton-electron repulsion, the exciton cannot populate the sites occupied by electrons (indicated by the crossed-out dashed path). The remaining available sites form an emergent excitonic lattice. (b) WC at various $\nu_{e}$ in zero-twist WSe2/WS2 [14, 28]. The remaining empty sites provide a lattice for exciton tunneling shown in panel (a).
@@
Two-band model. — We consider the lowest conduction (CB) and the highest valence (VB) bands of two TMD monolayers. Each band is labeled by the valley pseudospin, which is locked to the real spin degrees of freedom for energy scales lower than the corresponding spin-orbit splittings [34]. Stacking these layers with a tunable twist angle or (for heterobilayers) lattice mismatch generally leads to interlayer coupling with an enlarged spatial periodicity $a_{M}$. We specifically consider heterobilayers [35], as the band offsets therein allow us to capture the effect of interlayer tunneling by emergent superlattice potentials, which split the original electronic dispersion into moiré bands  [^3] Such band misalignment also suppresses the topological effects favored by layer-degenerate states [78], which is not the main focus of our work.. In addition, we consider doping electrons into the first conduction moiré band and focus on the regime with dilute electron-hole pairs [6, 37], generated via weak optical excitation. Note that within such a dilute regime, the emission properties are captured by the single excitation subspace, in which the spectrum of a master equation is mathematically equivalent to the eigenspectrum of a non-Hermitian Hamiltonian [1, 2, 3]. While our formalism can be straightforwardly generalized, for simplicity, all (doped and optically excited) electrons are assumed to be in the same layer.
@@
Zero doping. — At $\nu_{e}=0$, optical excitations from Eq. (1) are given by excitons. To capture the lowest composite particle, we employ the tight-binding approximation to the two-particle Hamiltonian operator from $\hat{\mathcal{H}}$  [^8] In the tight-binding regime, both the center-of-mass spatial fluctuations of the electron-hole pair from the moiré potential minima and the electron-hole separation are much smaller than the moiré period., whose eigenfunctions in the tight-binding limit can be approximated by moiré-Wannier orbitals [38]. The lowest energy orbital at each supersite $\bm{R}$, denoted as $w_{\bm{R}}(\bm{s}_{c},\bm{s}_{v})$ with $\bm{s}_{c}$ and $\bm{s}_{v}$ being the coordinates of the electron and hole, respectively, defines the corresponding exciton creation operator:
@@
Projecting $\hat{\mathcal{H}}$ onto these single-exciton basis states, the superlattice Hamiltonian becomes  [^9] Here, we drop the dipole-dipole interaction from the small out-of-plane transition dipole (not the permanent dipole) of interlayer excitons [51], which can be further suppressed via hybridization with intralayer excitons [80].:
@@
Finite doping with charge order. — We can generalize Eq. (4) to finite $\nu_{e}$, where doped electrons form WCs  [^11] The collective excitonic properties for hole-doped cases are qualitatively similar to electron-doped ones.. These charge orders could emerge from a generalized Hubbard model [53, 54], which naturally appears by projecting Eq. (1) onto the first moiré band of doped charges [38]. Importantly, since the exciton satisfies the tight-binding criterion, its hole remains bound to its original electron, preventing recombination with doped charges [55]. Further assuming that the electronic states being stable against the dynamics of dilute optical excitations, we can treat the WC as a spatially periodic detuning, consistent with experimental signatures [32, 33]. This detuning is characterized by the gap between $\omega_{\mathrm{ex}}$ and the energy of lowest three-body state (one doped electron together with the excited electron-hole pair) within a supercell, denoted as $\omega_{t}$. In particular, in the regime $|\omega_{t}-\omega_{\mathrm{ex}}|\gg|t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}|$, the low-energy optical excitations are either within the $\sim\omega_{\mathrm{ex}}$ or $\sim\omega_{t}$ manifold, depending on the sign of $\omega_{t}-\omega_{\mathrm{ex}}$. We focus specifically on the case $\omega_{\mathrm{ex}}\ll\omega_{t}$, which is realizable by interlayer excitons [38, 20], to study the collective behavior of these emitters (three-body states with $\omega_{\mathrm{ex}}\gg\omega_{t}$ correspond to trions [56], which is not the focus of this work). In this subspace, these excitons are still described by the Hamiltonian Eq. (4121212Note that the energy integrals are different for the excitations in the $\sim\omega_{\mathrm{ex}}$ and $\sim\omega_{t}$ manifolds, due to their distinction in wavefunctions, except that now $\{\bm{R}\}$ lies in the emergent lattices, which are complementary lattices of the WCs. Accordingly, different emergent lattices can be realized simply by accessing distinct WCs in moiré TMD bilayers, which can be tuned by the electron filling fraction $\nu_{e}$ (doped electrons per supercell) [53, 14, 28]. We list a few possible emergent lattices observed in WSe2/WS2 in Table 1.
@@
Figure 2: Collective excitonic lineshifts $\Delta_{\bm{Q}}$ (solid lines) and linewidths $\Gamma_{\bm{Q}}$ (dashed) emerging from charge-ordered zero-twist WSe2/WS2 with electron fillings (a) $\nu_{e}=0$ and (b) $\nu_{e}=\frac{1}{3}$. The vertical axes are displayed in units of $\gamma N_{\nu_{e}}$, with $N_{\nu_{e}}$ defined in Eq. (6). The horizontal axes represent the Bloch momentum $\bm{Q}$, which follows a piecewise-linear path through high-symmetry points in the Brillouin zone, as indicated by the hexagon in the inset of (a). Momenta within the light cone are indicated by the green shaded area (size enlarged for clarity). Different colors label distinct single-particle exciton bands. The parameters used are: $a_{M}=8.25$nm, $\omega_{\mathrm{ex}}=1.55$eV, and $a_{W}=2$nm [38, 26].
@@
Collective bands. — We proceed to study the eigenspectrum of Eq. (4). To capture the relevant physics, we set $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}=0$ in Eq. (5) throughout this work  [^13] Neglecting $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}$ is a valid approximation at large $a_{M}$ because it decays exponentially with $|\bm{R}-\bm{R}|/a_{M}$, whereas the dipole-dipole interaction only scales as a power law.. Diagonalization of the Hamiltonian Eq. (4) yields the emitter spectrum $\omega_{\mathrm{ex}}+\Delta_{\bm{Q}}-\frac{i}{2}\Gamma_{\bm{Q}}$ characterized by the center-of-mass Bloch momentum $\bm{Q}$. Note that, if the excitonic states at all $\bm{R}$ constructively interfere, $\Gamma_{\bm{Q}}$ should scale with $\gamma N_{\nu_{e}}(\theta)$ [1], where $N_{\nu_{e}}(\theta)$ is the ratio between the squared exciton wavelength and the unit cell area of the effective excitonic lattice at twisting angle $\theta$:
$N_{\nu_{e}}(\theta)=(1+\theta^{2}/\delta^{2})N_{\nu_{e}},\quad N_{\nu_{e}}=\lambda_{\mathrm{ex}}^{2}/\mathcal{A}_{\nu_{e}}.$ (6)
@@
Here $\mathcal{A}_{\nu_{e}}$ denotes the emergent unit cell area at electron filling $\nu_{e}$ and zero twist and $\delta$ defines the lattice mismatch between the two monolayers [26]. We present the collective spectrum in units of $\gamma N_{\nu_{e}}(\theta)$ hereafter to indicate the extent of constructive interference of the exciton eigenstates.
@@
Aside from these common properties, collective bands exhibit several distinct features at different $\nu_{e}$. We specifically illustrate how $\Delta_{0}$ and $\Gamma_{0}$ vary with $\nu_{e}$ in Fig. 3(a), as these quantities are directly accessible via reflection measurements at normal incidence [1]. First, different emergent lattices generally yield distinct numbers of bright levels, which serves as the most direct signature to identify certain WCs. For instance, two bright lines with equal and different $\Gamma_{0}$ can be associated to rectangular ($\nu_{e}=\frac{1}{2}$) and Kagome ($\nu_{e}=\frac{1}{4}$) lattices, respectively. In contrast, triangular and honeycomb lattices cannot be differentiated by bright-level counting.
Another optical signature of the WCs is the linewidth $\Gamma_{0}$ of the lowest state at each $\nu_{e}$, which is typically the fastest radiative rate. By comparing their values (which overall decreases with $\nu_{e}$), it is possible to determine whether the corresponding WCs are formed in experiments  [^14] While a similar comparison can be made for $\Delta_{0}$, we anticipate that additional factors beyond the scope of this study, including long-range exciton-electron interactions, may also influence the exciton energy. In spectral measurements, isolating $\Delta_{0}$ from those factors could be challenging..
@@
Finally, these spectral properties can also be tuned via twisting angle $\theta$. Fig. 3(b) illustrates the dependence of $\Gamma_{0}$ on $N_{\nu_{e}}(\theta)$. Crucially, for the same type of lattice geometry, the collective linewidths scale inversely with the emergent unit cell area, which can be confirmed by the analytic expression of the Green’s tensor at $\bm{Q}=0$. These scalings provide an indirect probe for the emergent lattice geometry: for instance, all triangular arrays (e.g., $\nu_{e}=0,\frac{2}{3}$) belong to the same line in Fig. 3(b). Therefore, the intercept in this logarithmic plot serves as another fingerprint of the WCs.
Berry curvature. — Topological features can also emerge in these collective excitonic bands. To see this, we introduce a magnetic field to break time-reversal symmetry and compute the total Berry curvature within the LC, $C_{\mathrm{LC}}$, for (directly) optically accessible states, as summarized in Table 1. Excitonic bands arising from triangular, honeycomb, and Kagome effective lattices yield $C_{\mathrm{LC}}=\pm 1$, whereas those from rectangular lattices remain topologically trivial. This behavior originates from the phase winding of the light-mediated coupling between valley doublets, which effectively manifests only in lattices with $C_{3}$ rotational symmetry, rather than from topological electronic bands [61].
Finally, we propose a far-field reflection experiment to detect $C_{\mathrm{LC}}$, where the scattering matrix element extracted from cross-circularly polarized input and output fields, $S_{-+}$, is measured. Specifically, its momentum dependence, controlled by incidence direction, possesses difference number of nodes for excitonic bands with different topological features. We leave the details to the End Matter.
Outlook. — Our formalism can be further generalized to describe the collective behavior of moiré excitons involving other strong correlations to study their interplay [62]. For instance, going beyond the single excitation subspace [63], nonlinearities inherent in these emitters can lead to a variety of interesting physical phenomena, including Dicke superradiance [64, 65, 66], optical bistability [67], leaky condensation [68], and phase space filling [69, 70, 71]. Another intriguing problem is whether the spin correlations that emerge near half-filling $\nu_{e}=1$, such as the magnetic polaron effect [72, 73] and kinetic magnetism [56], could play a role in the cooperative excitonic properties.
Acknowledgements. — We acknowledge A. Srivastava, S. Yelin and A. Asenjo-Garcia and D. Goncalves for useful discussions. The work at Maryland was supported by W911NF2010232, MURI FA9550-19-1-0399, DARPA HR00112530313 and Simons and Minta Martin Foundations. Y.-X.W. acknowledges support from a QuICS Hartree Postdoctoral Fellowship. Y.-Q. W. acknowledges the support from the JQI postdoctoral fellowship at the University of Maryland. D.E.C. acknowledges support from the European Union, under European Research Council grant agreement No 101002107 (NEWSPIN), FET-Open grant agreement No 899275 (DAALI) and EIC Pathfinder Grant No 101115420 (PANDA); the Government of Spain (Severo Ochoa Grant CEX2019-000910-S [MCIN/AEI/10.13039/501100011033]); QuantERA II project QuSiED, co-funded by the European Union Horizon 2020 research and innovation programme (No 101017733) and the Government of Spain (European Union NextGenerationEU/PRTR PCI2022-132945 funded by MCIN/AEI/10.13039/501100011033); Generalitat de Catalunya (CERCA program and AGAUR Project No. 2021 SGR 01442); Fundació Cellex, and Fundació Mir-Puig. All authors acknowledge the hospitality of the Kavli Institute for Theoretical Physics (KITP) supported by grant NSF PHY-1748958.
Appendix A End Matter